OpenStudy (darkbluechocobo):

Convert the function below to vertex form, showing your work for full credit. f(x) = x^2 + 2x + 3

OpenStudy (darkbluechocobo):

@ganeshie8

ganeshie8 (ganeshie8):

we use this formula \[\large a^2+2ab+b^2 = (a+b)^2\]

ganeshie8 (ganeshie8):

\(\large\begin{align} f(x) &= x^2 + 2x + 3\\& = x^2 + 2.1.x+3 \\& = x^2+2.1x + 1+2 \\&=(x+1)^2+2\end{align}\)

ganeshie8 (ganeshie8):

thats the vertex form

OpenStudy (darkbluechocobo):

@ganeshie8 sorry im back i have been sick so i had to go get medicine

ganeshie8 (ganeshie8):

Oh what happened? you should be resting and you can do homework later right

OpenStudy (darkbluechocobo):

aha nos i have to go somewhere this weekend :c so i gotta all of it done ahead of time